1.    Upon heating, NaN3(s) decomposes to Na(l) and N2(g). This reaction serves as a convenient laboratory preparation of pure nitrogen gas.

2 NaN3(s) ¾® 2 Na(l) + 3 N2(g)


a)       What number of moles of NaN3 is required to prepare 1.00 mol of N2?

 

 

 

b)      What mass of N2 is produced by the decomposition of 2.00 mol NaN3? (

 

 

 

c) What mass of Na is produced when 1.75 g of N2 is prepared? 


1
Expert's answer
2021-05-28T00:56:05-0400

We have given the reaction:


"2NaN_3 \rightarrow 2Na+3N_2"


a.) 3 moles of "N_2" gives 2 moles of "NaN_3"

1 mole of "N_2 \hspace{1mm} gives = \dfrac{2}{3 }" moles of "NaN_3"


b.) 2 moles of "NaN_3" gives 1 mole of "N_2"

1 mole of "NaN_3" give "= \dfrac{1}{3}" moles of "N_2 = \dfrac{1}{3} \times 28"


c.) Similarly we can also this part also.